# Imperative Learning: A Self-supervised Neural-Symbolic Learning Framework for Robot Autonomy

### v1

@@ 1 Introduction
While NeSy reasoning offers substantial potential, its specific application in the field of robotics is still in a nascent stage. A key reason for this is the emerging nature of the NeSy reasoning field itself, which has not yet reached a consensus on a rigorous definition (Kautz, 2022). On one side, a narrow definition views NeSy reasoning as an amalgamation of neural methods (data-driven) and symbolic methods, which utilize formal logic and symbols for knowledge representation and rule-based reasoning. Alternatively, the broader definition expands the scope of what constitutes a “symbol”. In this perspective, symbols are not only logical terms but also encompass any comprehensible, human-conceived concepts. This can include physical properties and semantic attributes, as seen in approaches involving concrete equations (Duruisseaux et al., 2023), logical programming (Delfosse et al., 2023), and programmable objectives (Yonetani et al., 2021a). This broad definition hence encapsulates reasoning related to physical principles, logical reasoning, geometrical analytics, etc.
@@
In this article, we explore the broader definition of NeSy reasoning for robot autonomy and introduce a novel self-supervised computational framework, which will be referred to as imperative learning (IL). It is designed to overcome three well-known issues of data-driven models for robot autonomy: (1) Generalization ability: Many data-driven systems, including reinforcement learning models (Banino et al., 2018; Albrecht et al., 2022), remain largely confined to their training environments, displaying limited generalization capabilities. One reason for this limitation is their inability to learn explicit commonsense rules, which hinders their effective transfer to new environments. This encourages us to delve into symbolic reasoning techniques. (2) Black-box nature: The underlying causal mechanisms of data-driven models are mostly unknown, while a problematic decision can have catastrophic side effects in robotic tasks (Dulac-Arnold et al., 2021). Consequently, these models frequently encounter difficulties when applied to real-world situations, which further encourages us to delve into NeSy systems. (3) Label intensiveness: Labeling data for robotic tasks like imitation learning (Zare et al., 2023) often incurs higher costs than tasks in computer vision due to the reliance on specialized equipment over basic human annotations (Ebadi et al., 2022). For example, annotating accurate ground truth for robot planning is extremely complicated due to the intricacies of robot dynamics. This underscores the critical need for efficient self-supervised learning methods. (3) Sub-optimality: Separately training neural and symbolic modules can result in sub-optimal integration due to compounded errors (Besold et al., 2021). This motivates the investigation of an end-to-end NeSy learning approach.
@@
  • We discuss several optimization strategies to tackle the technical challenges of IL. We present how to incorporate different optimization techniques into our IL framework including closed-form solution, first-order optimization, second-order optimization, constrained optimization, and combinatorial optimization.
  • To benefit the entire community, we demonstrated the effectiveness of IL for several tasks in robot autonomy including path planning, rule induction, optimal control, visual odometry, and multi-robot routing. We also released the source code to the public to inspire more robotics research using IL.
This article is inspired by and built on our previous works in different fields of robot autonomy including simultaneous localization and mapping (SLAM) (Fu et al., 2024), local planning (Yang et al., 2023), global planning (Chen et al., 2024), multi-agent routing (Guo et al., 2024), and feature matching (Zhan et al., 2024). These works introduced a prototype of IL in several different domains while failing to introduce a systematic methodology for NeSy learning in robot autonomy. This article fills this gap by formally defining IL, exploring various optimization challenges of IL in different robot autonomy tasks, and introducing new applications such as rule induction and optimal control. Additionally, we introduce the theoretical background for solving IL based on BLO, propose several practical solutions to IL by experimenting with the five distinct robot autonomy applications, and demonstrate the superiority of IL over state-of-the-art (SOTA) methods in their respective fields.
@@ 2.1 Bilevel Optimization
Bilevel optimization (BLO), first introduced by Bracken and McGill (1973), has been studied for decades. Classific approaches replaced the lower-level problem with its optimality conditions as constraints and reformulated the bilevel programming into a single-level constrained problem (Hansen et al., 1992; Gould et al., 2016; Shi et al., 2005; Sinha et al., 2017). More recently, gradient-based BLO has attracted significant attention due to its efficiency and effectiveness in modern machine learning and deep learning problems. Since this paper mainly focuses on the learning side, we will focus on gradient-based BLO methods in this section, and briefly discuss their limitations in robot learning problems.
@@
Despite these advancements, the application of BLO in robotics remains under-explored. This is particularly challenging in scenarios where convergence cannot be guaranteed, or where numerous constraints are impractical. For example, real-world problems are often highly non-convex, leading to optimization landscapes with many local minima and saddle points (Jadbabaie et al., 2019). Moreover, robotics problems frequently involve dynamic environments and real-time decision-making requirements that add further layers of difficulty to implementing BLO (Zhao et al., 2024a). We will explore the methods with assured convergence as well as those empirically validated by different tasks, especially from the perspective of robot autonomy.
@@ 3.2 Formulation
$\bm{\psi}=\operatorname*{\arg\min}_{\bm{\psi}\doteq[{\bm{\theta}},~{}{\bm{\gamma}}]}$ $U\left(f({\bm{\theta}},\bm{x}),g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*})\right),$ (1a)
s.t. $\bm{\phi}^{*}=\operatorname*{\arg\min}_{\bm{\phi}\doteq[{\bm{\mu}},~{}{\bm{\nu}}]}L(f({\bm{\theta}},\bm{x}),g({\bm{\mu}}),M({\bm{\gamma}},{\bm{\nu}})),$ (1b)
@@
where $\xi$ is a general constraint (either equality or inequality); $U$ and $L$ are the upper-level (UL) and lower-level (LL) cost functions; and $\bm{\psi}\doteq[{\bm{\theta}},{\bm{\gamma}}]$ are stacked UL variables and $\bm{\phi}\doteq[{\bm{\mu}},{\bm{\nu}}]$ are stacked LL variables, respectively. Alternatively, $U$ and $L$ are also referred to as the neural cost and symbolic cost, respectively. As aforementioned, the term “imperative” is used to denote the passive nature of the learning process: once optimized, the neural system $f$ in the UL cost will be driven to align with the LL reasoning engine $g$ (e.g., logical, physical, or geometrical reasoning process) with constraint $\xi$, so that it can learn to generate logically, physically, or geometrically feasible semantic attributes or predicates. In some application, $\bm{\psi}$ and $\bm{\phi}$ are also referred to as neuron-like and symbol-like parameters, respectively.
@@ 3.3.1 Unrolled Differentiation
$\bm{\Phi}_{t}(\bm{\phi}_{t-1};\bm{\psi})=\bm{\phi}_{t-1}-\eta\cdot\frac{\partial{L}(\bm{\phi}_{t-1},\bm{\bm{\psi}})}{\partial\bm{\phi}_{t-1}},$ (6)
@@ 3.3.2 Implicit Differentiation
In practice, the Hessian matrix can be too big to calculate and store, but we could bypass it by solving a linear system. Substitute (13) into (4), we have the UL gradient
@@ 4 Applications and Examples
4.1 Closed-form
@@
One possible problem in computing (3) is that the implicit components can contain expensive matrix inversions. Let us consider a simplified quadratic case in (1b) with ${\bm{\mu}}^{*}=\arg\min_{{\bm{\mu}}}\frac{1}{2}\|f({\bm{\theta}},x){\bm{\mu}}-y\|^{2}$. The closed-form solution takes the form of ${\bm{\mu}}^{*}=[f({\bm{\theta}},x)^{T}f({\bm{\theta}},x)]^{-1}y$, which can induce a computationally expensive inversion of a possibly large matrix $f({\bm{\theta}},x)^{T}f({\bm{\theta}},x)$. To address this problem, one can again formulate the matrix-inverse-vector computation $H^{-1}y$ as solving a linear system $\min_{v}\frac{1}{2}v^{T}Hv-v^{T}y$ using any optimization methods with efficient matrix-vector products.
@@
where the optimal path ${\bm{\mu}}^{*}$ is derived from the A reasoning $g$ given the node cost maps $M$. Thanks to its closed-form solution, this LL optimization is directly solvable. The UL optimization focuses on updating the network parameter ${\bm{\theta}}$ to generate the $h$ map. Given the impact of the $h$ map on the A’s search area, the UL cost $U$ is formulated as a combination of the search area $C_{a}({\bm{\mu}}^{*},{\bm{\nu}}^{*})$ and the path length $C_{l}({\bm{\mu}}^{*})$. This is mathematically represented as:
@@ 4.2 First-order Optimization
$\nabla_{\bm{\theta}}U=$ $\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}}+\bigg{[}\begin{array}[]{c}\frac{\partial U}{\partial{\bm{\mu}}^{*}}\\ \frac{\partial U}{\partial{\bm{\nu}}^{*}}\end{array}\bigg{]}^{T}\bigg{[}\begin{array}[]{c}\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\theta}}}\\ \frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\theta}}}\end{array}\bigg{]}$ (34)
$=$ $\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}}-\underbrace{\bigg{[}\begin{array}[]{c}\frac{\partial U}{\partial{\bm{\mu}}^{*}}\\ \frac{\partial U}{\partial{\bm{\nu}}^{*}}\end{array}\bigg{]}^{T}\bigg{[}\begin{array}[]{c}B,\;C\\ \widetilde{C},\;\widetilde{B}\end{array}\bigg{]}^{-1}}_{D}\bigg{[}\begin{array}[]{c}A\\ \widetilde{A}\end{array}\bigg{]}$ (41)
@@
$\min_{v}\quad\frac{1}{2}v^{T}\bigg{[}\begin{array}[]{c}B,\;C\\ \widetilde{C},\;\widetilde{B}\end{array}\bigg{]}v+\bigg{[}\begin{array}[]{c}\frac{\partial U}{\partial{\bm{\mu}}^{*}}\\ \frac{\partial U}{\partial{\bm{\nu}}^{*}}\end{array}\bigg{]}^{T}v.$ (42)
@@ 4.3 Constrained Optimization
Recall that $\phi=[{\bm{\mu}}^{T},{\bm{\nu}}^{T}]^{T}$ is a vector concatenating the symbol-like variables ${\bm{\mu}}$ and ${\bm{\nu}}$, and $\psi=[{\bm{\theta}}^{T},{\bm{\gamma}}^{T}]^{T}$ is a vector concatenating the neuron-like variables ${\bm{\theta}}$ and ${\bm{\gamma}}$. Therefore, the implicit gradient components can be expressed as differentiation of $\phi^{*}$ to $\psi$, and we have
$\nabla\phi^{*}(\psi)=\begin{pmatrix}\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\gamma}}}\\ \frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\gamma}}}\\ \end{pmatrix}.$ (46)
@@
$L_{\phi}=$ $\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\phi},\quad L_{\psi}=\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\psi},$ (48)
$L_{\phi\psi}=$ $\frac{\partial^{2}L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\psi\partial\phi}-\lambda\frac{\partial^{2}\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\psi\partial\phi},$
$H=$ $\frac{\partial^{2}L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial^{2}\phi}-\lambda\frac{\partial^{2}\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial^{2}\phi},$
and duel variable $\lambda\in\mathbb{R}$ satisfies $\lambda L_{\phi}=\frac{\partial L\xi(f,g({\bm{\mu}}),M({\bm{\gamma}},{\bm{\nu}}))}{\partial\phi}$.
@@
$\mathfrak{L}(\psi,\phi,\lambda)=L(f,g({\bm{\mu}}),M({\bm{\gamma}},{\bm{\nu}}))-\lambda\xi(M({\bm{\gamma}},{\bm{\nu}}),{\bm{\mu}},f).$ (49)
Assume $\phi^{*}$ to be the solution given input $\psi$. Then, a $\lambda$ exists such that the Lagrangian has zero gradients at $\phi^{*}$ and $\lambda$. Thus, taking derivatives w.r.t. $\phi$ and $\lambda$ yields
$\Big{[}\begin{array}[]{c}\frac{\partial L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\phi}-\lambda\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\phi}\\ \xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)\end{array}\Big{]}=\mathbf{0}.$ (50)
Note that if $\lambda=0$, then this means that the gradient $\frac{\partial L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\phi}=\mathbf{0}$ (i.e., the solution of the unconstrained problems already satisfies the constraint). If $\lambda>0$, then we can obtain from (50) that
$\frac{\partial L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\phi}=\lambda\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\phi},$ (51)
which means that the gradient of $L$ w.r.t. $\phi$ has the same direction as the gradient of constraint $\xi$ w.r.t. $\phi$ at the optimal solution ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$. Next, taking the derivative of (50) w.r.t. $\psi$ and noting that $\phi^{*}$ and $\lambda$ are functions of $\psi$, thus
$\frac{\partial^{2}L}{\partial\psi\partial\phi}+\frac{\partial^{2}L}{\partial^{2}\phi}\nabla\phi^{*}-\frac{\partial\xi}{\partial\phi}\nabla\lambda-\lambda(\frac{\partial^{2}\xi}{\partial\psi\partial\phi}+\frac{\partial^{2}\xi}{\partial^{2}\phi}\nabla\phi^{*})=\mathbf{0}$ (52a)
$\frac{\partial\xi}{\partial\psi}+\frac{\partial\xi}{\partial\phi}\nabla\phi^{*}=\mathbf{0},$ (52b)
@@
Because the inequality constraints are active at $\phi^{*}$ (i.e., $\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)=0$), the remaining proof exactly follows the same steps as in the equality case. ∎
While we have derived closed-form solutions for the equality constraint in (45) and the inequality constraint in (54), calculating the matrix inversions $H^{-1}$ and $[L_{\phi}H^{-1}L_{\phi}^{T}]^{-1}$ can be computationally expensive. To mitigate this, one may employ the approach discussed in Section 4.1, which involves approximating the matrix-inverse-vector product $H^{-1}y$ by solving the linear system $\min_{v}\frac{1}{2}v^{T}Hv-v^{T}y$. This complicated computing process can sometimes be simplified by specific robot settings. We next present an example of robot control illustrating this process.

### v2

@@ 1 Introduction
While NeSy reasoning offers substantial potential, its specific application in the field of robotics is still in a nascent stage. A key reason for this is the emerging nature of the NeSy reasoning field itself, which has not yet reached a consensus on a rigorous definition (Kautz, 2022). On one side, a narrow definition views NeSy reasoning as an amalgamation of neural methods (data-driven) and symbolic methods, which utilize formal logic and symbols for knowledge representation and rule-based reasoning. Alternatively, the broader definition expands the scope of what constitutes a “symbol”. In this perspective, symbols are not only logical terms but also encompass any comprehensible, human-conceived concepts. This can include physical properties and semantic attributes, as seen in approaches involving concrete equations (Duruisseaux et al., 2023), logical programming (Delfosse et al., 2023), and programmable objectives (Yonetani et al., 2021a). This broad definition hence encapsulates reasoning related to physical principles, logical reasoning, geometrical analytics, etc. In this context, many literatures exemplified NeSy systems, e.g., model-based reinforcement learning (Moerland et al., 2023), physics-informed networks (Karniadakis et al., 2021), and learning-aided tasks such as control (O’Connell et al., 2022), task scheduling (Gondhi and Gupta, 2017), and geometry analytics (Heidari and Iosifidis, 2024).
@@
In this article, we explore the broader definition of NeSy reasoning for robot autonomy and introduce a self-supervised NeSy learning framework, which will be referred to as imperative learning (IL). It is designed to overcome the well-known issues of existing learning frameworks for robot autonomy: (1) Generalization ability: Many data-driven systems, including reinforcement learning models (Banino et al., 2018; Albrecht et al., 2022), remain largely confined to their training environments, displaying limited generalization capabilities. One reason for this limitation is their inability to learn explicit commonsense rules, which hinders their effective transfer to new environments. This encourages us to delve into symbolic reasoning techniques. (2) Black-box nature: The underlying causal mechanisms of data-driven models are mostly unknown, while a problematic decision can have catastrophic side effects in robotic tasks (Dulac-Arnold et al., 2021). Consequently, these models frequently encounter difficulties when applied to real-world situations, which further encourages us to delve into NeSy systems. (3) Label intensiveness: Labeling data for robotic tasks like imitation learning (Zare et al., 2023) often incurs higher costs than tasks in computer vision due to the reliance on specialized equipment over basic human annotations (Ebadi et al., 2022). For example, annotating accurate ground truth for robot planning is extremely complicated due to the intricacies of robot dynamics. This underscores the critical need for efficient self-supervised learning methods. (4) Sub-optimality: Separately training neural and symbolic modules can result in sub-optimal integration due to compounded errors (Besold et al., 2021). This motivates the investigation of an end-to-end NeSy learning approach.
@@
  • We discuss several optimization strategies to tackle the technical challenges of IL. We present how to incorporate different optimization techniques into our IL framework including closed-form solution, first-order optimization, second-order optimization, constrained optimization, and discrete optimization.
  • To benefit the entire community, we demonstrated the effectiveness of IL for several tasks in robot autonomy including path planning, rule induction, optimal control, visual odometry, and multi-robot routing. We released the source code at https://sairlab.org/iseries/ to inspire more robotics research using IL.
This article is inspired by and built on our previous works in different fields of robot autonomy including local planning (Yang et al., 2023), global planning (Chen et al., 2024), simultaneous localization and mapping (SLAM) (Fu et al., 2024), feature matching (Zhan et al., 2024), and multi-agent routing (Guo et al., 2024). These works introduced a prototype of IL in several different domains while failing to introduce a systematic methodology for NeSy learning in robot autonomy. This article fills this gap by formally defining IL, exploring various optimization challenges of IL in different robot autonomy tasks, and introducing new applications such as rule induction and optimal control. Additionally, we introduce the theoretical background for solving IL based on BLO, propose several practical solutions to IL by experimenting with the five distinct robot autonomy applications, and demonstrate the superiority of IL over state-of-the-art (SOTA) methods in their respective fields.
@@ 2.1 Bilevel Optimization
Bilevel optimization (BLO), first introduced by Bracken and McGill (1973), has been studied for decades. Classific approaches replaced the lower-level problem with its optimality conditions as constraints and reformulated the bilevel programming into a single-level constrained problem (Hansen et al., 1992; Gould et al., 2016; Shi et al., 2005; Sinha et al., 2017). More recently, gradient-based BLO has attracted significant attention due to its efficiency and effectiveness in modern machine learning and deep learning problems. Since this paper mainly focuses on the learning side, we will concentrate on gradient-based BLO methods, and briefly discuss their limitations in robot learning problems.
@@
Despite these advancements, the application of BLO in robotics remains under-explored. This is particularly challenging in scenarios where convergence cannot be guaranteed, or where numerous constraints are impractical. For example, real-world problems are often highly non-convex, leading to optimization landscapes with many local minima and saddle points (Jadbabaie et al., 2019). Moreover, robotics problems frequently involve dynamic environments and real-time decision-making requirements that add further layers of difficulty to implementing BLO (Zhao et al., 2024a). We will explore the methods with assured convergence as well as those empirically validated by various tasks, especially from the perspective of robot autonomy.
@@ 3.2 Formulation
$\min_{\bm{\psi}\doteq[{\bm{\theta}}^{\top},~{}{\bm{\gamma}}^{\top}]^{\top}}$ $U\left(f({\bm{\theta}},\bm{x}),g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*})\right),$ (1a)
s.t. $\bm{\phi}^{*}\in\operatorname*{\arg\min}_{\bm{\phi}\doteq[{\bm{\mu}}^{\top},~{}{\bm{\nu}}^{\top}]^{\top}}L(f({\bm{\theta}},\bm{x}),g({\bm{\mu}}),M({\bm{\gamma}},{\bm{\nu}})),$ (1b)
@@
where $\xi$ is a general constraint (either equality or inequality); $U$ and $L$ are the upper-level (UL) and lower-level (LL) cost functions; and $\bm{\psi}\doteq[{\bm{\theta}}^{\top},{\bm{\gamma}}^{\top}]^{\top}$ are stacked UL variables and $\bm{\phi}\doteq[{\bm{\mu}}^{\top},{\bm{\nu}}^{\top}]^{\top}$ are stacked LL variables, respectively. Alternatively, $U$ and $L$ are also referred to as the neural cost and symbolic cost, respectively. As aforementioned, the term “imperative” is used to denote the passive nature of the learning process: once optimized, the neural system $f$ in the UL cost will be driven to align with the LL reasoning engine $g$ (e.g., logical, physical, or geometrical reasoning process) with constraint $\xi$, so that it can learn to generate logically, physically, or geometrically feasible semantic attributes or predicates. In some applications, $\bm{\psi}$ and $\bm{\phi}$ are also referred to as neuron-like and symbol-like parameters, respectively.
@@ 3.3.1 Unrolled Differentiation
$\bm{\Phi}_{t}(\bm{\phi}_{t-1};\bm{\psi})=\bm{\phi}_{t-1}-\eta\cdot\frac{\partial{L}(\bm{\phi}_{t-1};\bm{\bm{\psi}})}{\partial\bm{\phi}_{t-1}},$ (6)
@@ 3.3.2 Implicit Differentiation
In practice, the Hessian matrix can be too big to calculate and store [^2] For instance, assume both UL and LU costs have a network with merely 1 million ($10^{6}$) parameters ($32$-bit float numbers), thus each network only needs a space of $10^{6}\times 4\text{Byte}=4\mathrm{M}B$ to store, while their Hessian matrix needs a space of $(10^{6})^{2}\times 4\text{Byte}=4\mathrm{T}B$ to store. This indicates that a Hessian matrix cannot even be explicitly stored in the memory of a low-power computer, thus directly calculating its inversion is more impractical., but we could bypass it by solving a linear system. Substitute (13) into (4), we have the UL gradient
@@
4 Applications and Examples
4.1 Closed-form Solution
@@
One possible problem in computing (3) is that the implicit components can contain expensive matrix inversions. Let us consider a simplified quadratic case in (1b) with ${\bm{\mu}}^{*}=\arg\min_{{\bm{\mu}}}\frac{1}{2}\|f({\bm{\theta}},x){\bm{\mu}}-y\|^{2}$. The closed-form solution takes the form of ${\bm{\mu}}^{*}=[f({\bm{\theta}},x)^{\top}f({\bm{\theta}},x)]^{-1}y$, which can induce a computationally expensive inversion of a possibly large matrix $f({\bm{\theta}},x)^{\top}f({\bm{\theta}},x)$. To address this problem, one can again formulate the matrix-inverse-vector computation $H^{-1}y$ as solving a linear system $\min_{v}\frac{1}{2}v^{\top}Hv-v^{\top}y$ using any optimization methods with efficient matrix-vector products.
@@
where the optimal path ${\bm{\mu}}^{*}$ is derived from the A reasoning $g$ given the node cost maps $M$. Thanks to its closed-form solution, this LL optimization is directly solvable. The UL optimization focuses on updating the network parameter ${\bm{\theta}}$ to generate the $h$ map. Given the impact of the $h$ map on the search area of A, the UL cost $U$ is formulated as a combination of the search area $C_{a}({\bm{\mu}}^{*},{\bm{\nu}}^{*})$ and the path length $C_{l}({\bm{\mu}}^{*})$. This is mathematically represented as:
@@ 4.2 First-order Optimization
$\nabla_{\bm{\theta}}U=$ $\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}}+\bigg{[}\begin{array}[]{c}\frac{\partial U}{\partial{\bm{\mu}}^{*}}\\ \frac{\partial U}{\partial{\bm{\nu}}^{*}}\end{array}\bigg{]}^{\top}\bigg{[}\begin{array}[]{c}\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\theta}}}\\ \frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\theta}}}\end{array}\bigg{]}$ (34)
$=$ $\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}}-\underbrace{\bigg{[}\begin{array}[]{c}\frac{\partial U}{\partial{\bm{\mu}}^{*}}\\ \frac{\partial U}{\partial{\bm{\nu}}^{*}}\end{array}\bigg{]}^{\top}\bigg{[}\begin{array}[]{c}B,\;C\\ \widetilde{C},\;\widetilde{B}\end{array}\bigg{]}^{-1}}_{D}\bigg{[}\begin{array}[]{c}A\\ \widetilde{A}\end{array}\bigg{]}$ (41)
@@
$\min_{v}\quad\frac{1}{2}v^{\top}\bigg{[}\begin{array}[]{c}B,\;C\\ \widetilde{C},\;\widetilde{B}\end{array}\bigg{]}v+\bigg{[}\begin{array}[]{c}\frac{\partial U}{\partial{\bm{\mu}}^{*}}\\ \frac{\partial U}{\partial{\bm{\nu}}^{*}}\end{array}\bigg{]}^{\top}v.$ (42)
@@ 4.3 Constrained Optimization
Recall that $\bm{\phi}=[{\bm{\mu}}^{\top},{\bm{\nu}}^{\top}]^{\top}$ is a vector concatenating the symbol-like variables ${\bm{\mu}}$ and ${\bm{\nu}}$, and $\bm{\psi}=[{\bm{\theta}}^{\top},{\bm{\gamma}}^{\top}]^{\top}$ is a vector concatenating the neuron-like variables ${\bm{\theta}}$ and ${\bm{\gamma}}$. Therefore, the implicit gradient components can be expressed as differentiation of $\bm{\phi}^{*}$ to $\bm{\psi}$, and we have
$\nabla\bm{\phi}^{*}(\bm{\psi})=\begin{pmatrix}\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\gamma}}}\\ \frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\gamma}}}\\ \end{pmatrix}.$ (46)
@@
$L_{\bm{\phi}}=$ $\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial{\bm{\phi}}},\quad L_{\bm{\psi}}=\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial{\bm{\psi}}},$ (48)
$L_{\bm{\phi\psi}}=$ $\frac{\partial^{2}L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\bm{\psi}\partial\bm{\phi}}-\lambda\frac{\partial^{2}\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\bm{\psi}\partial\bm{\phi}},$
$H=$ $\frac{\partial^{2}L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial^{2}\bm{\phi}}-\lambda\frac{\partial^{2}\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial^{2}\bm{\phi}},$
and duel variable $\lambda\in\mathbb{R}$ satisfies $\lambda L_{\bm{\phi}}=\frac{\partial L\xi(f,g({\bm{\mu}}),M({\bm{\gamma}},{\bm{\nu}}))}{\partial\bm{\phi}}$.
@@
$\mathfrak{L}(\bm{\psi},\bm{\phi},\lambda)=L(f,g({\bm{\mu}}),M({\bm{\gamma}},{\bm{\nu}}))-\lambda\xi(M({\bm{\gamma}},{\bm{\nu}}),{\bm{\mu}},f).$ (49)
Assume $\bm{\phi}^{*}$ to be the solution given input $\bm{\psi}$. Then, a $\lambda$ exists such that the Lagrangian has zero gradients at $\bm{\phi}^{*}$ and $\lambda$. Thus, taking derivatives w.r.t. $\bm{\phi}$ and $\lambda$ yields
$\Bigg{[}\begin{array}[]{c}\frac{\partial L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\bm{\phi}}-\lambda\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\bm{\phi}}\\ \xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)\end{array}\Bigg{]}=\mathbf{0}.$ (50)
Note that if $\lambda=0$, then this means that the gradient $\frac{\partial L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\bm{\phi}}=\mathbf{0}$ (i.e., the solution of the unconstrained problems already satisfies the constraint). If $\lambda>0$, then we can obtain from (50) that
$\frac{\partial L(f,g({\bm{\mu}}^{*}),M({\bm{\gamma}},{\bm{\nu}}^{*}))}{\partial\bm{\phi}}=\lambda\frac{\partial\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)}{\partial\bm{\phi}},$ (51)
which means that the gradient of $L$ w.r.t. $\bm{\phi}$ has the same direction as the gradient of constraint $\xi$ w.r.t. $\bm{\phi}$ at the optimal solution ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$. Taking the derivative of (50) w.r.t. $\bm{\psi}$ and noting that $\bm{\phi}^{*}$ and $\lambda$ are functions of $\bm{\psi}$, thus
$\frac{\partial^{2}L}{\partial\bm{\psi}\partial\bm{\phi}}+\frac{\partial^{2}L}{\partial^{2}\bm{\phi}}\nabla\bm{\phi}^{*}-\frac{\partial\xi}{\partial\bm{\phi}}\nabla\lambda-\lambda(\frac{\partial^{2}\xi}{\partial\bm{\psi}\partial\bm{\phi}}+\frac{\partial^{2}\xi}{\partial^{2}\bm{\phi}}\nabla\bm{\phi}^{*})=\mathbf{0},$ (52a)
$\frac{\partial\xi}{\partial\bm{\psi}}+\frac{\partial\xi}{\partial\bm{\phi}}\nabla\bm{\phi}^{*}=\mathbf{0},$ (52b)
@@
Because the inequality constraints are active at $\bm{\phi}^{*}$ (i.e., $\xi(M({\bm{\gamma}},{\bm{\nu}}^{*}),{\bm{\mu}}^{*},f)=0$), the remaining proof exactly follows the same steps as in the equality case. ∎
While we have derived closed-form solutions for the equality constraint in (45) and the inequality constraint in (54), calculating the matrix inversions $H^{-1}$ and $[L_{\bm{\phi}}H^{-1}L_{\bm{\phi}}^{\top}]^{-1}$ can be computationally expensive. To mitigate this, one may employ the approach discussed in Section 4.1, which involves approximating the matrix-inverse-vector product $H^{-1}y$ by solving the linear system $\min_{v}\frac{1}{2}v^{\top}Hv-v^{\top}y$. This complicated computing process can sometimes be simplified by specific robot settings. We next present an example of robot control illustrating this process.